OpenStudy (anonymous):

Please help!! Test on this tomorrow! If f varies directly as g and inversely as the square of h, and f=20 when g=50 and h=5, find f when g=18 and h=6. I totally understand what direct and inverse mean and how to set up the problem. I just don't know what to do whenever it's varied to something SQUARED. Do I square the number given? I really need to see what's done step by step and then I'll be confident I'm getting the other ones like this correct. Thanks!

OpenStudy (anonymous):

@satellite73 @geekfromthefutur @paki @sourwing @agent0smith

OpenStudy (anonymous):

My teacher explained the y=kx and the y=k/x and all that, just not what to do when it's squared. Never seen anything similar in my book so I want to make sure I'm on the right track

OpenStudy (paki):

what you want to ask...?

OpenStudy (anonymous):

How do I solve a problem like this? I understand to PLUG IN the numbers to their proper places, but (for instance) should I SQUARE h? And If I do square h, when I find k and plug it back in, should I square the new h value again?

OpenStudy (paki):

hmmm... ok... for squaring the number.... just put square over them and solve... other steps are similar like solving the direct and indirect proportions....

OpenStudy (anonymous):

So I set up, f=gk/h 20=(5^2)k/50 so k = 100 plug back in now f=18(100)/6 f=300 correct answer?

OpenStudy (paki):

look, what is the value of g and h... check it in youir formula...

OpenStudy (anonymous):

g=50 and h=5 in the first, then g=18 and h=6

OpenStudy (anonymous):

My bad, I didn't square the second time. 6 squared is 36. so it would be f = 50

OpenStudy (paki):

so how you put the values in,,, f=gk/h

OpenStudy (paki):

hmmm :-)

OpenStudy (anonymous):

Simple fact error lol my fault. Thank you for your help though I understand it a lot better now! so f=50 is the correct answer!

OpenStudy (paki):

yeah ;-) and pleasure